# Unleashing the Power of Emojis in Texts via Self-supervised Graph Pre-Training

### v1

@@ Abstract
Emojis have gained immense popularity on social platforms, serving as a common means to supplement or replace text. However, existing data mining approaches generally either completely ignore or simply treat emojis as ordinary Unicode characters, which may limit the model’s ability to grasp the rich semantic information in emojis and the interaction between emojis and texts. Thus, it is necessary to release the power of emojis in social media data mining. To this end, we first construct a heterogeneous graph consisting of three types of nodes, i.e., post, word and emoji nodes to improve the representation of different elements in posts. The edges are also well-defined to model how these three elements interact with each other. To facilitate the sharing of information among post, word and emoji nodes, we propose a graph pre-training framework for text and emoji co-modeling, which contains two graph pre-training tasks: node-level graph contrastive learning and edge-level link reconstruction learning. Extensive experiments on the Xiaohongshu and Twitter datasets with two types of downstream tasks demonstrate that our approach proves significant improvement over previous strong baseline methods.
@@ 1 Introduction
  • We conduct extensive experiments on datasets from different social media platforms (i.e., Xiaohongshu and Twitter) to demonstrate that our approach achieves better performance on various downstream tasks (e.g. popularity prediction and sentiment prediction) than strong baselines. Our model outperforms the baseline model by 2% - 10%. Besides, our model can be further used for the emoji generation task.
@@ 3 Methodology
In this section, we first construct the heterogeneous graph with three types of nodes: posts, emojis and words, to capture the information and interactions among them (§ 3.1). Then, based on the heterogeneous graph, we design a graph pre-training framework to model the comprehensive interaction among all types of nodes (§ 3.2). After that, we show how to use our proposed method in various downstream tasks, i.e., popularity prediction, emoji generation and sentiment classification(§ 3.3). Figure 1 illustrates the overview of our pre-training and fine-tuning stages.
@@ 3.3 Downstream Tasks
To evaluate our proposed framework, we apply our emoji embeddings to two downstream tasks: popularity prediction and sentiment analysis. For all downstream tasks, the task input is a social-media post $P=\{T_{p};E_{p}\}$, where $T_{p}=\{w_{p_{1}},w_{p_{2}},...w_{p_{m}}\}$ is the text in the post, $w_{p_{i}}$ means the $i$-th word in the post text $T_{p}$, and $m$ indicates the number of words in $T_{p}$. $E_{p}=\{e_{p_{1}},e_{p_{2}},...,e_{p_{n}}\}$ denotes the emojis in $P$, $e_{j}$ is the $j$-th emoji, and $n$ represents the number of emojis in $E_{p}$. The representations of each word $w_{p_{i}}$ and emoji $e_{p_{j}}$ can be initialized to $h_{w_{pi}}\in\mathbb{R}^{d}$ and $h_{e_{pj}}\in\mathbb{R}^{d}$ via our graph pre-training framework.***If a word or an emoji (in a downstream post) does not appear during pre-training, it will be discarded to avoid the out-of-vocabulary problem. We first sample a sub-graph of this node from the heterogeneous graph via a random walk starting from the node, and then input the sub-graph into our GNN encoder model to get its representation, which also serves as the embedding of the node. This embedding is then used for downstream tasks.
@@ 4 Experiment
In this section, we first introduce the implementation details of our proposed pre-training framework (§ 4.1), and then evaluate our framework on two downstream tasks, i.e., popularity prediction (§ 4.2) and sentiment classification (§ 4.3). Also we use illustrative exampleS on emoji generation (§ 4.4) for further demonstration. Finally, we conduct ablation studies to show the effectiveness of each pre-training task (§ 4.5).
@@ 4.2 Popularity Prediction
Dataset Our first dataset mainly comes from the Chinese social media platform Xiaohongshuhttps://www.xiaohongshu.com/. We collect over 4 million Xiaohongshu posts through web crawling, and all the information are desensitized. In addition to the content of the post, other information mainly includes the author’s Xiaohongshu ID, post title (text information), number of likes, number of collects, number of comments, and publishing time. The user information mainly includes the number of followers, the total number of likes, and whether the account is an official certified account. Note that this Xiaohongshu dataset will not be public due to ethical considerations (e.g., the post might contain individuals’ privacy). Our popularity label is derived from the “liked count”, “collected count”, and “comments count” fields in the Xiaohongshu dataset. Given the continuous nature of this data, to simplify it into a categorical task, we cluster them based on the distribution trend of popularity according to the data percentile of [0-50), [50-80), and [80-100], resulting in three categories (low, medium, and high). Thus, the ratio of category labels is low:medium:high = 5:3:2, indicating the imbalanced labels in the data. Given this, we use the weighted labels to mitigate the imbalance issue. The weights of the labels are the inverses of their counts, as described in Eq. 8. The statistic of dataset is shown in table 4.
@@
  • Preprocessing methods. We use the traditional emoji preprocessing methods introduced in section 2 as baselines: (1) Remove emojis: we use vectors of zeroes as the embedding of emojis to eliminate the information contained in them. (2) Keep emojis: we use vectors of numbers randomly generated from Gaussian distribution as the embedding of emojis, meaning that they are different tokens but the model has no prior information on emojis. We generated two versions of embeddings to avoid randomness, which are denoted as Keep V1 and Keep V2. (3) Translate emojis: we translate emojis into their descriptions (e.g., replace “ [Uncaptioned image] ” with the word “car”), and then use our backbone LLMs to encode emojis. In this way, emojis are treated just as normal words.
  • Emoji representation learning method. These baseline model utilize the Emoji2vec (Eisner et al., 2016) and the Emoji Co-occurence (Illendula and Yedulla, 2018) methods to obtain emojis’ embedding. For those emojis not included in the model, we initiate them to vectors of zeroes.
@@
As illustrated in Table 1, our method outperforms all four baselines on all three sub-types of popularity prediction tasks. This finding suggests that emojis offer extra information for popularity prediction by comparing our model with the remove baseline. Besides, traditional methods generally fail or have a minor effect on extracting information from emojis. When the emoji embedding is initiated by Gaussian random noise, the F1 score even decreases. This is because no improvement in the prediction performance can be attributed to the absence of semantic information in Gaussian noise. The translate baseline utilizes LLMs to extract emoji information by translating emojis into words and then encoding them. Different from Gaussian noise, emoji tokens contain valuable semantics. However, due to there is only a very small part of emojis can be translated, we can infer that incorporating emoji information results in nearly no improvement in the F1 score. Emoji2vec enhances the extraction of emoji information and achieves better performance compared to translate. Our method shows the highest F1 score among all models, indicating its strong capacity for extracting information from emojis. Detailed precision and recall results are shown in appendix B. To further demonstrate the superior performance of our model over baseline models on downstream tasks, we conducted two-sample t-tests between our model and each baseline model. The result shows that our model significantly outperforms baseline models at the 0.05 significance level. Detailed results of the hypothesis tests are included in appendix C.
@@
4.3 Sentiment Classification
@@
4.4 Illustrative Example on Emoji Generation
Figure 3: The figure shows an illustrative example of how our model can be combined with existing LLMs to generate suitable emojis for posts to achieve high popularity.
Figure 4: The emoji generation case outputs by GPT-4o. The left part is generated by GPT only. The right part is by entering the output of our model in the prompt.
To further demonstrate that our pre-training framework can effectively learn the relationship between posts, words, and emojis, we also show an illustrative example in a real-world scenario: post emoji generation. Specifically, this scenario receives post $P$ that only contains text $T_{p}$ as a input, and inserts emojis into the post. The generation process includes three steps: (1) we use an emoji prediction to determine the most $n_{e}$ relevant emojis $E_{r}=\{e_{r_{1}},e_{r_{2}},...,e_{r_{n}}|e_{r_{i}}\in\mathcal{E}\}$ with the post $P$. The similarity between each emoji $e_{i}$ and the post $P$ is calculated as $W_{ep}(h_{e_{i}}\oplus h_{t})^{T}$, where $W_{ep}$ is trainable parameters. (2) Then, word prediction will find the most matched $n_{w}$ words from $T_{p}$ for each emoji in $E_{r}$. The similarity between emoji $e_{r_{i}}$ and word $w_{p_{j}}$ is calculated as $W_{ew}(h_{e_{r_{i}}}\oplus h_{w_{p_{j}}})^{T}$. To train $W_{ep}$ and $W_{ew}$, it can be regarded as a binary classification for every emoji (matched or unmatched with posts and words). (3) In the last step, we turn $E_{r}$ and the corresponding matched words into a prompt for GPT-4o model to output a post with emojis.
Results & Analyses We show several cases to demonstrate the efficiency of our model. At each time, we enter the original post without emojis plus the emojis and their correlated words generated by our model to form a prompt for LLMs models. The results are shown in Figure 4. Compared to the output post generated only by prompting GPT to insert emojis, entering the results ($E_{r}$ and the corresponding matched words) of our model helps both in the richness of emojis and the accuracy of the position to insert emojis (more cases are given in Appendix D).
@@ 4.4 Illustrative Example on Emoji Generation
4.5 Ablation Study
The node-level sub-graph contrastive learning task is separately conducted on different types of nodes, and thus includes three sub-tasks, i.e., post-node, word-node and emoji-node contrastive learning. The edge-level link reconstruction also consists of three sub-tasks, i.e., post-emoji, word-post and emoji-word link reconstruction. To evaluate the effectiveness of these sub-tasks, we conduct ablations by removing each of them in our pre-training framework, and evaluate the model performance on popularity prediction (number of likes).
As shown in Table 3, the performance of our original framework exceeds every other variant model, indicating that our model gain information from every part of pre-training tasks. In conclusion, the ablation study shows that Emoji representation learning can benefit from all the pre-training sub-tasks, leading to enhanced prediction performance.
@@ Acknowledgements
Appendix A Dataset Description Statistic
Metric Mean Std Min .25 .50 .75 Max
Liked Count 799.0 5198.5 0.0 66.0 148.0 334.0 357163.0
Collected Count 468.3 3431.5 0.0 25.0 75.0 189.0 281984.0
Comments Count 46.3 140.7 0.0 9.0 22.0 46.0 8652.0
Emoji Count/Post 9.3 8.4 2.0 4.0 7.0 12.0 316.0
Post Length 391.0 258.2 13.0 191.0 335.0 535.0 18500.0
Table 4: Dataset Description Statistic
Appendix B Full Results of Popularity Prediction
@@
Appendix C Significance Test
We conduct a two-sample t-test between our proposed model and baselines. Experimental results are shown in Table 7, which indicate that our model is significantly better than baselines with t-test p < 0.05 in each test.
Backbone BERT GPT2 RoBERTa OPT-1.3B LlAMA2-7B
F1 score (Weighted) - Prediction of # of likes
v.s. Remove 0.0013 0.0000 0.0000 0.0000 0.0000
v.s. Keep V1 0.0344 0.0023 0.0000 0.0000 0.0000
v.s. Keep V2 0.0000 0.0015 0.0000 0.0002 0.0000
v.s. Translate 0.0000 0.0003 0.0000 0.0001 0.0000
v.s. Emoji2vec 0.0462 0.0000 0.0000 0.0000 0.0000
F1 score (Weighted) - Prediction of # of comments
v.s. Remove 0.0416 0.0000 0.0000 0.0000 0.0014
v.s. Keep V1 0.0000 0.0000 0.0000 0.0000 0.0000
v.s. Keep V2 0.0000 0.0000 0.0000 0.0000 0.0000
v.s. Translate 0.0000 0.0000 0.0000 0.0000 0.0000
v.s. Emoji2vec 0.0693 0.0000 0.0000 0.0000 0.0014
F1 score (Weighted) - Prediction of # of collects
v.s. Remove 0.0257 0.0000 0.0000 0.0087 0.0052
v.s. Keep V1 0.0000 0.0000 0.0000 0.0000 0.0000
v.s. Keep V2 0.0000 0.0000 0.0000 0.0000 0.0000
v.s. Translate 0.0000 0.0000 0.0000 0.0000 0.0000
v.s. Emoji2vec 0.0482 0.0000 0.0000 0.0329 0.0020
Table 7: Results of t-test. Each score represents the average p-value for t-test
Appendix D More Emoji Generation Showcases
We show more cases of the output generated by GPT-3.5-turbo collaborated with our framework in Figure 5. It is illustrated that by using the output of our model, the emoji generation task witnesses great improvement.
Figure 5: Three cases for the emoji generation task. The left part is the output by using GPT-3.5-turbo directly, while the right part is the output generated by GPT-3.5-turbo collaborated with our framework.
Appendix E Bad Cases Study
In this bad case study of sentiment analysis task, which are shown in Table 8, we observe discrepancies between the model’s predictions and the actual sentiments. These instances highlight several critical challenges.
Text Prediction Ground Truth
Table 8: Some bad cases in sentiments analysis

### v2

@@ Abstract
Emojis have gained immense popularity on social platforms, serving as a common means to supplement or replace text. However, existing data mining approaches generally either completely ignore or simply treat emojis as ordinary Unicode characters, which may limit the model’s ability to grasp the rich semantic information in emojis and the interaction between emojis and texts. Thus, it is necessary to release the power of emojis in social media data mining. To this end, we first construct a heterogeneous graph consisting of three types of nodes, i.e., post, word and emoji nodes to improve the representation of different elements in posts. The edges are also well-defined to model how these three elements interact with each other. To facilitate the sharing of information among post, word and emoji nodes, we propose a graph pre-training framework for text and emoji co-modeling, which contains two graph pre-training tasks: node-level graph contrastive learning and edge-level link reconstruction learning. Extensive experiments on the Xiaohongshu and Twitter datasets with two types of downstream tasks demonstrate that our approach proves significant improvement over previous strong baseline methods. [^1] https://github.com/ginkoeric/Self-supervised-Graph-Pre-training-for-Emoji
@@ 1 Introduction
  • We conduct extensive experiments on datasets from different social media platforms (i.e., Xiaohongshu and Twitter) to demonstrate that our approach achieves better performance on various downstream tasks (e.g. popularity prediction and sentiment prediction) than strong baselines. Our model outperforms the baseline model by 2% - 10%.
@@ 3 Methodology
In this section, we first construct the heterogeneous graph with three types of nodes: posts, emojis and words, to capture the information and interactions among them (§ 3.1). Then, based on the heterogeneous graph, we design a graph pre-training framework to model the comprehensive interaction among all types of nodes (§ 3.2). After that, we show how to use our proposed method in various downstream tasks, i.e., popularity prediction and sentiment classification(§ 3.3). Figure 1 illustrates the overview of our pre-training and fine-tuning stages.
@@ 3.3 Downstream Tasks
To evaluate our proposed framework, we apply our emoji embeddings to two downstream tasks: popularity prediction and sentiment analysis. For all downstream tasks, the task input is a social-media post $P=\{T_{p};E_{p}\}$, where $T_{p}=\{w_{p_{1}},w_{p_{2}},...w_{p_{m}}\}$ is the text in the post, $w_{p_{i}}$ means the $i$-th word in the post text $T_{p}$, and $m$ indicates the number of words in $T_{p}$. $E_{p}=\{e_{p_{1}},e_{p_{2}},...,e_{p_{n}}\}$ denotes the emojis in $P$, $e_{j}$ is the $j$-th emoji, and $n$ represents the number of emojis in $E_{p}$. The representations of each word $w_{p_{i}}$ and emoji $e_{p_{j}}$ can be initialized to $h_{w_{pi}}\in\mathbb{R}^{d}$ and $h_{e_{pj}}\in\mathbb{R}^{d}$ via our graph pre-training framework. [^2] If a word or an emoji (in a downstream post) does not appear during pre-training, it will be discarded to avoid the out-of-vocabulary problem. We first sample a sub-graph of this node from the heterogeneous graph via a random walk starting from the node, and then input the sub-graph into our GNN encoder model to get its representation, which also serves as the embedding of the node. This embedding is then used for downstream tasks.
@@ 4 Experiment
In this section, we first introduce the implementation details of our proposed pre-training framework (§ 4.1), and then evaluate our framework on two downstream tasks, i.e., popularity prediction (§ 4.2) and sentiment classification (§ 4.3).
Metric Mean Std Min .25 .50 .75 Max
Liked Count 799.0 5198.5 0.0 66.0 148.0 334.0 357163.0
Collected Count 468.3 3431.5 0.0 25.0 75.0 189.0 281984.0
Comments Count 46.3 140.7 0.0 9.0 22.0 46.0 8652.0
Emoji Count/Post 9.3 8.4 2.0 4.0 7.0 12.0 316.0
Post Length 391.0 258.2 13.0 191.0 335.0 535.0 18500.0
Table 1: Dataset Description Statistic.
@@ 4.2 Popularity Prediction
Dataset Our first dataset mainly comes from the Chinese social media platform Xiaohongshu [^3] https://www.xiaohongshu.com/. We collect over 4 million Xiaohongshu posts through web crawling, and all the information are desensitized. In addition to the content of the post, other information mainly includes the author’s Xiaohongshu ID, post title (text information), number of likes, number of collects, number of comments, and publishing time. The user information mainly includes the number of followers, the total number of likes, and whether the account is an official certified account. Note that this Xiaohongshu dataset will not be public due to ethical considerations (e.g., the post might contain individuals’ privacy). Our popularity label is derived from the “liked count”, “collected count”, and “comments count” fields in the Xiaohongshu dataset. Given the continuous nature of this data, to simplify it into a categorical task, we cluster them based on the distribution trend of popularity according to the data percentile of [0-50), [50-80), and [80-100], resulting in three categories (low, medium, and high). Thus, the ratio of category labels is low:medium:high = 5:3:2, indicating the imbalanced labels in the data. Given this, we use the weighted labels to mitigate the imbalance issue. The weights of the labels are the inverses of their counts, as described in Eq. 8. The statistic of dataset is shown in table 1.
@@
  • Preprocessing methods: We use the traditional emoji preprocessing methods introduced in section 2 as baselines: (1) Remove emojis: we use vectors of zeroes as the embedding of emojis to eliminate the information contained in them. (2) Keep emojis: we use vectors of numbers randomly generated from Gaussian distribution as the embedding of emojis, meaning that they are different tokens but the model has no prior information on emojis. We generated two versions of embeddings to avoid randomness, which are denoted as Keep V1 and Keep V2. (3) Translate emojis: we translate emojis into their descriptions (e.g., replace “ [Uncaptioned image] ” with the word “car”), and then use our backbone LLMs to encode emojis. In this way, emojis are treated just as normal words.
  • Emoji representation learning method: These baseline model utilize the Emoji2vec (Eisner et al., 2016) and the Emoji Co-occurence (Illendula and Yedulla, 2018) methods to obtain emojis’ embedding. For those emojis not included in the model, we initiate them to vectors of zeroes.
@@
As illustrated in Table 2, our method outperforms all four baselines on all three sub-types of popularity prediction tasks. This finding suggests that emojis offer extra information for popularity prediction by comparing our model with the remove baseline. Besides, traditional methods generally fail or have a minor effect on extracting information from emojis. When the emoji embedding is initiated by Gaussian random noise, the F1 score even decreases. This is because no improvement in the prediction performance can be attributed to the absence of semantic information in Gaussian noise. The translate baseline utilizes LLMs to extract emoji information by translating emojis into words and then encoding them. Different from Gaussian noise, emoji tokens contain valuable semantics. However, due to there is only a very small part of emojis can be translated, we can infer that incorporating emoji information results in nearly no improvement in the F1 score. Emoji2vec enhances the extraction of emoji information and achieves better performance compared to translate. Our method shows the highest F1 score among all models, indicating its strong capacity for extracting information from emojis. Detailed precision and recall results are shown in appendix A.
To further demonstrate the superior performance of our model over baseline models on downstream tasks, we conducted two-sample t-tests between our model and each baseline model. The result shows that our model significantly outperforms baseline models at the 0.05 significance level. Detailed results of the hypothesis tests are shown in Table 3.
@@
Backbone BERT GPT2 RoBERTa OPT-1.3B LlAMA2-7B
F1 score (Weighted) - Prediction of # of likes
v.s. Remove 0.0013 0.0000 0.0000 0.0000 0.0000
v.s. Keep V1 0.0344 0.0023 0.0000 0.0000 0.0000
v.s. Keep V2 0.0000 0.0015 0.0000 0.0002 0.0000
v.s. Translate 0.0000 0.0003 0.0000 0.0001 0.0000
v.s. Emoji2vec 0.0462 0.0000 0.0000 0.0000 0.0000
F1 score (Weighted) - Prediction of # of comments
v.s. Remove 0.0416 0.0000 0.0000 0.0000 0.0014
v.s. Keep V1 0.0000 0.0000 0.0000 0.0000 0.0000
v.s. Keep V2 0.0000 0.0000 0.0000 0.0000 0.0000
v.s. Translate 0.0000 0.0000 0.0000 0.0000 0.0000
v.s. Emoji2vec 0.0693 0.0000 0.0000 0.0000 0.0014
F1 score (Weighted) - Prediction of # of collects
v.s. Remove 0.0257 0.0000 0.0000 0.0087 0.0052
v.s. Keep V1 0.0000 0.0000 0.0000 0.0000 0.0000
v.s. Keep V2 0.0000 0.0000 0.0000 0.0000 0.0000
v.s. Translate 0.0000 0.0000 0.0000 0.0000 0.0000
v.s. Emoji2vec 0.0482 0.0000 0.0000 0.0329 0.0020
Table 3: Results of t-test. Each score represents the average p-value for t-test.
@@
Text Prediction Ground Truth
Table 5: Some bad cases in sentiments analysis.
@@
Ablation Study The node-level sub-graph contrastive learning task is separately conducted on different types of nodes, and thus includes three sub-tasks, i.e., post-node, word-node and emoji-node contrastive learning. The edge-level link reconstruction also consists of three sub-tasks, i.e., post-emoji, word-post and emoji-word link reconstruction.
To evaluate the effectiveness of these sub-tasks, we conduct ablations by removing each of them in our pre-training framework, and evaluate the model performance on popularity prediction (number of likes). As shown in Table 6, the performance of our original framework exceeds every other variant model, indicating that our model gain information from every part of pre-training tasks. In conclusion, the ablation study shows that Emoji representation learning can benefit from all the pre-training sub-tasks, leading to enhanced prediction performance.
4.3 Sentiment Classification
Bad Cases Study In this bad case study of sentiment analysis task, which are shown in Table 5, we observe discrepancies between the model’s predictions and the actual sentiments. These instances highlight several critical challenges.
@@ Acknowledgements
Appendix A Full Results of Popularity Prediction